How many unit cells can be made out of an 18 carat pure 197 g sample of gold (fcc structure, A= 197 u)
 

Grade:12th pass

2 Answers

RiQ
6 years ago
24k=100%gold22k=91.7%gold18k=75%goldHere 75%of 197g is pure gold.0.75moles of gold .No. of atoms (Au) = 0.75×6×10²³ = 45×10²².1 FCC str have 4 atoms .Hence no. of unit cells = 11.25×10²¹ .U can visit RiQ UR Classes for more help on you tube . Thanks !
RiQ
6 years ago
24k=100%gold22k=91.7%gold18k=75%goldHere 75%of 197g is pure gold.0.75moles of gold .No. of atoms (Au) = 0.75×6×10²³ = 45×10²².1 FCC str have 4 atoms .Hence no. of unit cells = 11.25×10²¹ .U can visit RiQ UR Classes for more help on you tube . Thanks !.